OpenStudy (anonymous):

Simplify the expression. 22 • 23 A. 45 B. 46 C. 26 D. 25

OpenStudy (compassionate):

Twenty-two times twenty-three is neither of these.

OpenStudy (lovelyanna):

Is that dot for multiplication? If so, none of the multiple choice answers are correct. Its 506.

OpenStudy (anonymous):

If your doing addition, the answer would be A. :P

OpenStudy (anonymous):

@Nial @LovelyAnna @Compasionate its suposed to be like this oops \[2^{2} • 2^{3}\] \[A. 4^{5} \] \[B. 4^{6}\] \[C. 2^{6}\] \[D. 2^{5}\]

OpenStudy (anonymous):

@Niall

OpenStudy (anonymous):

@Compassionate

OpenStudy (lovelyanna):

Makes much more sense! \[2^2*2^3=4*8=32\] So 2^5=32. Should be D.

OpenStudy (anonymous):

The answer is D. :)

OpenStudy (lovelyanna):

@Niall, thanks!

OpenStudy (anonymous):

Yw! I always give medals to best answers :P

OpenStudy (anonymous):

Thank you

OpenStudy (lovelyanna):

Lol, Thanks! :)

OpenStudy (lovelyanna):

No prob.